OpenStudy (anonymous):

How do you solve this simultaneous equation? 2x + 3y = 13 2xy + 5y^2 - 4x^2 = 41

OpenStudy (anonymous):

Use substitution: y = (13 - 2x)/3 and put it into your second equation. Hopefully you will get a quadratic in x which you can solve (hopefully b^2 - 4ac >= 0). There will be two values for x, each corresponding to a different value of y, unless you have a repeated x, in which case the y will be distinct.

OpenStudy (anonymous):

yeah. you could use the first equation for value of y and x, it's optional. then substitute value of y or x to the second equation to find x or y

OpenStudy (anonymous):

I've tried doing that but i can't seem to get the answer which is x=2 or -8/1/2. y=3 or 10. ;/

OpenStudy (anonymous):

I'll do it in a few mins after I've had my breakfast if it isn't done it by then...

OpenStudy (anonymous):

Okay, thank youu! (:

OpenStudy (anonymous):

Right...this is going to take a while to type just to give you the very gory last details...

OpenStudy (anonymous):

haha okay :)

OpenStudy (anonymous):

Rearrange 2x+3y=13 to give: y=(13-2x)/3. Plug this into your second equation to get: \[\frac{2x(13-2x)}{3} + \frac{5(13-2x)^2}{9} - 4x^2 - 41 = 0\] now multiply the whole thing by 3x9 = 27 to remove the fractions: \[18x(13-2x) + 15(169 - 52x + 4x^2) - 108x^2 - 1107 = 0\] next isolate the coefficients so we can write our familiar quadratic equation: \[x^2(60-36-108) + x(234-780) + (2535 - 1107) = 0\] now just calculate: \[-84x^2 - 546x + 1428 = 0\] simplify this by dividing by -6: \[ 14x^2 + 91x - 238 = 0\] Now use the quadratic formula because it's a lovely thing :p \[x = \frac{-91 \pm \sqrt{91^2 + 4(14)(238)}}{28} = \frac{-91 \pm \sqrt{21609}}{28} = \frac{-91 \pm 147}{28}\] so the positive one is x = 2 and the negative one is x=-8.5. Now we use the equation from before: \[y = \frac{13 - 2x}{3}\] in order to find two y-values, one associated with x=2, the other associated with x=-8.5. Substituting x=2 into y=(13-2x)/3 gives y = 3. Substituting x=-8.5 into y=(13-2x)/3 gives y = 10. So (x,y) solutions are: (2,3) and (-8.5, 10).

OpenStudy (anonymous):

Hope this is clear!

OpenStudy (anonymous):

you can pay me in medals :D

OpenStudy (anonymous):

thank you so much :) i understand now